# Certifiably Optimal Anisotropic Rotation Averaging

### v1

@@ 4.3 A stronger convex relaxation
can be computed while explicitly forcing $Q_{ij}\in SO(3)$. The minimum value of $\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ is 00 if $\mathbf{\Upsilon}-\mathbf{N}\succeq 0$ and $\infty$- ∞ otherwise. Let $\mathcal{I}_{SO(3)}$ be the indicator function,
@@
The last term can be identified as $-\mathcal{I}^{*}_{SO(3)}(\Upsilon_{ij})$ where $\mathcal{I}^{*}_{SO(3)}$ is the convex conjugate [34] of $\mathcal{I}_{SO(3)}$. We thus obtain the dual problem
@@
The constraint $X_{ij}\in\operatorname{conv}(SO(3))$ has been shown to be equivalent to a semidefinite constraint [38, 37]. A $3\times 3$ matrix $Y$ is in $\operatorname{conv}(SO(3))$ if and only if
$\mathcal{A}(Y)+{\mathtt{I}}\succeq 0,$ (27)
where $\mathcal{A}(Y)=$
{+pmatrix}$[colsep=0pt,rowsep=0.5pt]-Y_{11}-Y_{22}+Y_{33}&Y_{13}+Y_{31}Y_{12}-Y_{21}Y_{23}+Y_{32}\\ Y_{13}+Y_{31}Y_{11}-Y_{22}-Y_{33}Y_{23}-Y_{32}Y_{12}+Y_{21}\\ Y_{12}-Y_{21}Y_{23}-Y_{32}Y_{11}+Y_{22}+Y_{33}Y_{31}-Y_{13}\\ Y_{23}+Y_{32}Y_{12}+Y_{21}Y_{31}-Y_{13}-Y_{11}+Y_{22}-Y_{33}.$
@@ 5 Experiments
We implemented the SDP program using the conic splitting solver [31] with the JuMP [24] wrapper in Julia [^4] Code can be found at: https://ylochman.github.io/anisotropic-ra. The cost matrix in (SDP-cSO(3)) is down-scaled by the average (across the observed relative poses in the scene) of the largest eigenvalue of the computed Hessian matrix, which proved beneficial empirically. The absolute and relative feasibilities are set to $10^{-5}$ and $10^{-6}$, respectively, and infeasibility tolerance is $10^{-8}$. The number of iterations is limited to 500 000500000500\,000500 000. In practice, the methods converge in far fewer iterations to meet the stopping criterion. While it is likely that dedicated solvers are much more efficient, we remark that generalization of efficient methods such as [14] may not be straightforward since they rely heavily on properties $O(d)$, which are not sufficient to ensure $SO(3)$ membership.
@@
We compare [^5] Complementing results on both synthetic and real data can be found in the supplementary material. three approaches: 1) SDP-O(3)-iso is the standard rotation averaging approach that uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, 2) SDP-O(3)-aniso uses the anisotropic objective function but ignores the $\operatorname{conv}(SO(3))$ constraints, and 3) SDP-cSO(3) is the proposed method that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints. For completeness of the synthetic study, we also include SDP-cSO(3)-iso, which uses the isotropic objective and the $\operatorname{conv}(SO(3))$ constraints. The reported rotation error wrt. ground truth is the difference $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where $R_{i}$ are the ground truth absolute rotations and $R_{i}^{*}$ are the estimated ones. Note that to remove gauge freedom, we align the two sets of rotations $\{R_{i}\}_{i}$ and $\{R_{i}^{*}\}_{i}$, by applying a global rotation $V$ to all rotations in $\{R_{i}\}_{i}$ such that the rotation error achieves its minimum wrt. $V$.
@@ 5.1 Synthetic experiments
In Figure 3 we evaluate the effect of using the proposed anisotropic error measurements versus regular chordal distances together with the proposed relaxation versus the standard one. We vary the number of cameras from $2$ to $100$, and the proportion of missing data from $0\%$ to $90\%$. For each configuration pair, we run $100$ instances of the problem. In the top row of Figure 3, we report the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$. As expected, using the proposed method (i.e., both new objective and new relaxation) consistently leads to lower errors. We record the time-to-solution of the solver, shown in the bottom row of Figure 3. In this synthetic setting, the proposed relaxation is more efficient. Our intuition is that the introduction of local constraints—$X_{ij}\in\text{conv}(SO(3))$—reduces the oscillations in the employed first-order SDP solver and therefore leads to faster time-to-solution for optimizing both isotropic and anisotropic costs with the proposed constraints.
@@ 5.2 Real experiments
[…]
We report the obtained rank for each method and the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where the ground truth $\{R_{i}\}_{i}$ is from the final reconstruction obtained using the pipeline of [29]. When SDP-O(3)-aniso does not give a solution $X^{*}$ of rank 3, we compute the closest rank 3 approximation to $X^{*}$ by setting all but the three largest eigenvalues to 00. We then compute a factorization $X^{*}=VV^{\top}$, where $V$ has $3$ columns. From each consecutive $3\times 3$ block $V_{i}$ of $V$ we then compute the closest rotation matrix which we take to be the approximate solution $R_{i}^{*}$.
@@
LU Sphinx (70 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0944 2
@@
Round Church (92 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1399 6
@@
UWO (114 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3142 14
@@
Tsar Nikolai I (89 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1170 7
@@
Vercingetorix (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3146 2
@@
Eglise Du Dome (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0546 4
@@
King’s College (77 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1656 4
@@
Kronan (131 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2149 18
@@
Alcatraz (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1763 19
@@
Museum Barcelona (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2255 22
@@
Temple Singapore (157 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2646 31
@@
Table 1: Results on real data from [15, 29]. SDP-O(3)-iso uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-O(3)-aniso uses the proposed anisotropic objective but ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-cSO(3) is the proposed approach that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints.
@@ 6 Conclusions
In this work, we propose a way to incorporate anisotropic costs in a certifiably optimal rotation averaging framework. We demonstrate how traditional solvers designed to optimize chordal distances fail to optimize the new objective due to ignoring the determinant constraint. Our new SDP formulation enforces $\operatorname{conv}(SO(3))$ and our empirical evaluation shows that this relaxation is able to recover global minima. We evaluated the new approach on both synthetic and real datasets and obtained more accurate reconstructions than isotropic methods in most cases, confirming recent observations based on local optimization. We believe that this work will provide a good entry for research into anisotropic rotation averaging, a subarea of rotation averaging that has mostly been set aside because of the focus on isotropic cases thus far.
This work largely focuses on the modeling side and understanding the shortcomings of existing relaxations. We aim to leverage standard SDP solvers; however, the increased strength of our SDP formulation often comes at the cost of more expensive optimization. Designing a dedicated efficient algorithm is an important future direction.
@@ Appendix A First-Order Uncertainty Propagation
We leverage Laplace’s approximation, which allows us to identify the Hessian of a non-linear least-squares minimization task as the precision matrix of the posterior $p(\tilde{\bm{\omega}}_{ij}|\text{image $i$ and $j$})$ (up to a positive scale). In the following we apply first-order uncertainty propagation to move from the minimal axis-angle parametrization $\bm{\omega}\in\mathbb{R}^{3}$ to a direct parametrization of rotation matrices $R_{ij}\in\mathbb{R}^{3\times 3}$.
@@
$\mathtt{M}^{-1}=\Sigma_{[\Delta\bm{\omega}]_{\times}}=\mathtt{J}\Sigma_{\Delta\bm{\omega}}\mathtt{J}^{\top}=\mathtt{J}\mathtt{H}^{-1}\mathtt{J}^{\top}.$ (A.1)
@@
is linear in $R$ and matches $\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}$ to first order. We consider
@@
in accordance with Section 4.1. We relate $[\Delta\bm{\omega}]_{\times}$ and $R$ via $[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}}$ and therefore obtain
@@
A seemingly different solution $\mathtt{M}^{\prime}$ can be obtained by factoring $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})$ differently,
@@ Appendix B Parameterization of $R$
There are different ways of parametrizing rotations. In the main paper we use the $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ which lead to the objective of the form $-\langle M\tilde{R},R\rangle$. In this section, we investigate the effects of switching to other possible parameterizations. In particular, we start by letting $R=e^{[\bm{\omega}]_{\times}}=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$ and look at its approximations.
@@
The natural choices for $\alpha$ are 00, $1/2$ and $1$. Setting $\alpha=0$ yields the formulation employed in the main paper. We found that for the other values of $\alpha$ the solution of anisotropic rotation averaging is the same. However, when using the Hessian matrix computed from the Jacobian of the initial parameterization, $R=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$, the results vary when optimizing the anisotropic cost for different values of $\alpha$. In our synthetic experiments the best solution is often obtained with $\alpha=1/2$, however, this setting does not outperform the proposed formulation overall.
Appendix C Experimental Details
@@
C.1 Synthetic experiments
Figure C.1: The configuration of the cameras (top left). Rotation error wrt. ground truth $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$ (top right, bottom) for the increasing noise scale around one of the axes.
We set up a synthetic graph with three cameras as shown in Figure C.1. The estimated relative rotations corresponding to the black dashed edges are “certain”, i.e. $\Delta\bm{\omega}_{ij}$ are drawn from $\mathcal{N}(0,\varepsilon{\mathtt{I}})$, where $\varepsilon=0.001$. The estimated relative rotation of the gray edge has varying uncertainty around one of the three axes, e.g., for the x- (red) axis, the noise covariance is $\operatorname{diag}(\sigma,\varepsilon,\varepsilon)$, where $\sigma$ varies from $0.01$ to $0.3$. The effects on the error are shown in Figure C.1. The proposed method relies on certain relative rotations and gives an accurate solution, while the standard isotropic approach is negatively affected by the single noisy relative rotation.

### v2

@@ 4.3 A stronger convex relaxation
can be computed while explicitly forcing $Q_{ij}\in SO(3)$. The minimum value of $\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ is 0 if $\mathbf{\Upsilon}-\mathbf{N}\succeq 0$ and -$\infty$ otherwise. Let $\mathcal{I}_{SO(3)}$ be the indicator function,
@@
The last term can be identified as $\mathcal{I}^{*}_{SO(3)}(\Upsilon_{ij})$ where $\mathcal{I}^{*}_{SO(3)}$ is the convex conjugate [37] of $\mathcal{I}_{SO(3)}$. We thus obtain the dual problem
@@
The constraint $X_{ij}\in\operatorname{conv}(SO(3))$ has been shown to be equivalent to a semidefinite constraint [41, 40]. A $3\times 3$ matrix $Y$ is in $\operatorname{conv}(SO(3))$ if and only if $\mathcal{A}(Y)+{\mathtt{I}}\succeq 0$, where $\mathcal{A}(Y)=$
$\tiny\+pmatrix[colsep=0pt,rowsep=0.5pt]-Y_{11}-Y_{22}+Y_{33}&Y_{13}+Y_{31}&Y_{12}-Y_{21}&Y_{23}+Y_{32}\\ Y_{13}+Y_{31}&Y_{11}-Y_{22}-Y_{33}&Y_{23}-Y_{32}&Y_{12}+Y_{21}\\ Y_{12}-Y_{21}&Y_{23}-Y_{32}&Y_{11}+Y_{22}+Y_{33}&Y_{31}-Y_{13}\\ Y_{23}+Y_{32}&Y_{12}+Y_{21}&Y_{31}-Y_{13}&-Y_{11}+Y_{22}-Y_{33}.$
@@ 5 Experiments
We implemented the SDP program using the conic splitting solver [34] with the JuMP [27] wrapper in Julia [^6] The code is available at: https://github.com/ylochman/anisotropic-ra. The cost matrix in (SDP-cSO(3)) is down-scaled by the average (across the observed relative poses in the scene) of the largest eigenvalue of the computed Hessian matrix, which proved beneficial empirically. The absolute and relative feasibilities are set to $10^{-5}$ and $10^{-6}$, respectively, and infeasibility tolerance is $10^{-8}$. The number of iterations is limited to 500 000500\,000. In practice, the methods converge in far fewer iterations to meet the stopping criterion. While it is likely that dedicated solvers are much more efficient, we remark that generalization of efficient methods such as [16] may not be straightforward since they rely heavily on properties $O(d)$, which are not sufficient to ensure $SO(3)$ membership.
@@
We compare [^7] Complementing results on synthetic and real data can be found in the supplementary material. These also include comparison to the anisotropic extension of the spectral approach [2]. three approaches: 1) SDP-O(3)-iso is the standard rotation averaging approach that uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, 2) SDP-O(3)-aniso uses the anisotropic objective function but ignores the $\operatorname{conv}(SO(3))$ constraints, and 3) SDP-cSO(3) is the proposed method that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints. For completeness of the synthetic study, we also include SDP-cSO(3)-iso, which uses the isotropic objective and the $\operatorname{conv}(SO(3))$ constraints. The reported rotation error wrt. ground truth is the difference $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where $R_{i}$ are the ground truth absolute rotations and $R_{i}^{*}$ are the estimated ones. Note that to remove gauge freedom, we align the two sets of rotations $\{R_{i}\}_{i}$ and $\{R_{i}^{*}\}_{i}$, by applying a global rotation $V$ to all rotations in $\{R_{i}\}_{i}$ such that the rotation error achieves its minimum wrt. $V$.
@@ 5.1 Synthetic experiments
[…]
@@ 5.2 Real experiments
We present the results of applying the proposed framework to a number of public structure from motion datasets  [^8] Available at:
https://www.maths.lth.se/matematiklth/personal/calle/dataset/dataset.html
from [17, 32]. For every pair of images (with 5 or more point matches), we use RANSAC and the 5 point solver [31] to generate an initial solution to the pairwise geometry which is refined, using bundle adjustment [45], into a local minimum of the sum-of-squared reprojection errors. Here we fix gauge freedom by assuming that the first camera is fixed, the second camera is at distance $1$ from the first, and represent 3D points with homogeneous coordinates with norm $1$. We then compute the second order Taylor approximation of the objective function and marginalize over the position of the second camera and 3D points. This gives us a quadratic function, in the axis-angle variable that represents the relative rotation, as described in Section 4.1 and Figure 1, that we use for the anisotropic objective.
In Table 1 we report the obtained rank for each method and the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where the ground truth $\{R_{i}\}_{i}$ is from the final reconstruction obtained using the pipeline of [32]. When SDP-O(3)-aniso does not give a solution $X^{*}$ of rank 3, we compute the closest rank 3 approximation to $X^{*}$ by setting all but the three largest eigenvalues to 0. We then compute a factorization $X^{*}=VV^{\top}$, where $V$ has $3$ columns. From each consecutive $3\times 3$ block $V_{i}$ of $V$ we then compute the closest rotation matrix which we take to be the approximate solution $R_{i}^{*}$.
@@
LU Sphinx 70 cameras 85% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.0944 2
@@
Round Church 92 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1399 6
@@
UWO 114 cameras 77% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.3142 14
@@
Tsar Nikolai I 89 cameras 87% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1170 7
@@
Vercingetorix 69 cameras 77% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.3146 2
@@
Eglise Du Dome 69 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.0546 4
@@
King’s College 77 cameras 100% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1656 4
@@
Kronan 131 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2149 18
@@
Alcatraz 133 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1763 19
@@
Mus. Barcelona 133 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2255 22
@@
Temple Singapore 157 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2646 31
@@
Table 1: Results on real data from [17, 32]. SDP-O(3)-iso uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-O(3)-aniso uses the proposed anisotropic objective but ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-cSO(3) is the proposed approach that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints. The first column provides with the information about the dataset: the number of cameras and the percentage of indefinite matrices $M_{ij}$.
@@ 6 Conclusions
In this work, we propose to incorporate anisotropic costs in a certifiably optimal rotation averaging framework. We demonstrate how existing solvers designed for chordal distances fail to provide sensible solutions for the new objective as these ignore the determinant constraint. Our new SDP formulation enforces $\operatorname{conv}(SO(3))$, and our empirical evaluation shows that this relaxation is able to recover global minima. We evaluated the new approach on both synthetic and real datasets and obtained more accurate reconstructions than isotropic methods in most cases, confirming recent observations based on local optimization. We believe that our work serves as a good entry point for further investigating the anisotropic rotation averaging setting. This work focuses mainly on modeling aspects and understanding the shortcomings of existing relaxations. We aim to leverage standard SDP solvers; however, the increased strength of our SDP formulation often comes at the cost of more expensive optimization. Designing a dedicated efficient algorithm is an important future direction.
@@ Appendix A First-Order Uncertainty Propagation
[…]
@@
$\Sigma_{[\Delta\bm{\omega}]_{\times}}=\mathtt{J}\Sigma_{\Delta\bm{\omega}}\mathtt{J}^{\top}=\mathtt{J}\mathtt{H}^{-1}\mathtt{J}^{\top}.$ (A.1)
@@
is linear in $R$ and matches $\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}$ to first order for a suitable matrix $\mathtt{M}$ to be determined. We consider
@@
in accordance with Section 4 of the main paper. We relate $\Delta\bm{\omega}]_{\times}$ and $R$ via $\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}}$ and therefore obtain
@@
Since $\Delta\bm{\omega}]_{\times}$ is (skew-)symmetric, a seemingly different solution $\mathtt{M}^{\prime}$ can be obtained by factoring $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})$ differently,
@@ Appendix B Parameterization of $R$
There are different ways of parametrizing rotations. In the main paper we use the $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ which lead to the objective of the form $\langle M\tilde{R},R\rangle$. In this section, we investigate the effects of switching to other possible parameterizations. In particular, we start by letting $R=e^{[\bm{\omega}]_{\times}}=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$ and look at its approximations.
@@
The natural choices for $\alpha$ are 0, $1/2$ and $1$. Setting $\alpha=0$ yields the formulation employed in the main paper. We found that for the other values of $\alpha$ the solution of anisotropic rotation averaging is the same. However, when using the Hessian matrix computed from the Jacobian of the initial parameterization, $R=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$, the results vary when optimizing the anisotropic cost for different values of $\alpha$. In our synthetic experiments the best solution is often obtained with $\alpha=1/2$, however, this setting does not outperform the proposed formulation overall.
Appendix C Spectral method for anisotropic costs
We recall that our main objective is
$\sum_{i,j}\left\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\right\rangle.$ (C.13)
Let us assume the noisefree setting, i.e. $\tilde{R}_{ij}=R_{i}R_{j}^{\top}$. The cost matrix $\mathtt{N}$ is then given by
$\mathtt{N}=\begin{pmatrix}&\vdots&\\ \cdots&M_{ij}R_{j}R_{i}^{\top}&\cdots\\ &\vdots&\end{pmatrix}$. (C.14)
Consequently,
$\begin{split}\mathtt{N}\mathtt{R}&=\begin{pmatrix}\vdots\\ \sum_{i}M_{ij}R_{j}R_{i}^{\top}R_{i}\\ \vdots\end{pmatrix}=\begin{pmatrix}\vdots\\ \sum_{i}M_{ij}R_{j}\\ \vdots\end{pmatrix}\\ &=\underbrace{\begin{pmatrix}\ddots&&\\ &\sum_{i}M_{ij}&\\ &&\ddots\end{pmatrix}}_{=:\mathtt{D}}\mathtt{R}.\end{split}$ (C.15)
Hence, $\mathtt{N}\mathtt{R}=\mathtt{D}\mathtt{R}$ or
$\mathtt{D}^{-1}\mathtt{N}\mathtt{R}=\mathtt{R},$ (C.16)
and $\mathtt{R}$ can be extracted as the eigenspace corresponding to the eigenvalue of 1. Since $\mathtt{N}$ does not necessarily have rank-3 (even in the noise-free case), we obtain $\mathtt{R}$ as the right singular vectors corresponding to the three smallest singular values of $\mathtt{N}-{\mathtt{I}}$. We include results for this method in Table D.1 introduced in the next section.
Appendix D Experimental Details
@@
D.1 Synthetic experiments
Figure D.1: The configuration of the cameras (top left). Rotation error wrt. ground truth $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$ (top right, bottom) for the increasing noise scale around one of the axes.
We set up a synthetic graph with three cameras as shown in Figure D.1. The estimated relative rotations corresponding to the black dashed edges are “certain”, i.e. $\Delta\bm{\omega}_{ij}$ are drawn from $\mathcal{N}(0,\varepsilon{\mathtt{I}})$, where $\varepsilon=0.001$. The estimated relative rotation of the gray edge has varying uncertainty around one of the three axes, e.g., for the x- (red) axis, the noise covariance is $\operatorname{diag}(\sigma,\varepsilon,\varepsilon)$, where $\sigma$ varies from $0.01$ to $0.3$. The effects on the error are shown in Figure D.1. The proposed method relies on certain relative rotations and gives an accurate solution, while the standard isotropic approach is negatively affected by the single noisy relative rotation.